A sawmill makes boards that are used in construction projects. The board lengths of two-by-fours produced by this sawmill are uniformly distributed from 6 ft to 12 ft.
Question 1: Describe the graph of this distribution.
Question 2: Determine the probability that a board selected at random will be at least 10 ft long.
1
Expert's answer
2020-07-20T18:07:11-0400

Solution:

Q1: The probability density function and comutative distribution function on the interval [6;12] are P(x) and D(x) respectively.

"P(x) = \begin{cases}
 0 &\text{if }x<6 \\
 1/6 &\text{if } 6\le x \le 12 \\ 0 &\text{if }x > 12 
\end{cases}"


The graph P(x) is piecewise constant.





"D(x) = \begin{cases}
 0 &\text{if }x<6 \\
 (x-6)/6 &\text{if } 6\le x \le 12 \\ 1 &\text{if }x > 12 
\end{cases}"


The graph D(x) is piecewise linear.





Q2: "P(X\ge 10) =D(+\infty)-D(10)=1-(10-6)/6=1/3"